OpenStudy (anonymous):

(7-8i)^2 multiply, answer in a+bi form

OpenStudy (lalaly):

(7-8i)(7-8i)= 49-56i-56i-64= -15 -112i

jimthompson5910 (jim_thompson5910):

Answer is \[\large (7-8i)^2=-15-112i\] See attached for how I got that answer.